Friday November 8 Preble Company manufacturos one product. Its variable manufacturing overhead is applied to production basedFriday November 8 2. What is the materials quantity variance for March?(ndicate the effect of each variance by selecting F4. If Preble had purchased 186.000 pounds of materials at $6.80 per pound and used 175.000 pounds in production, what would bFriday November 8 7. What is the direct labor efficiency variance for March2(Indicate the effect of each variance by seloctin10. What is the variable overhead efficioncy variance for March?(Indicate the offect of each variance by Friday November 8 seFriday November 8 13. What is the spending variance related to advertising? (Indicate the effect of each variance by selectin
Answer #1

Solution 1:

Raw materials cost would be included in the company’s flexible budget for March = 34000*5*$7 = $1,190,000

Solution 2:

Materials quantity variance for March = (SQ - AQ) * SP = (34000*5 - 175000) * $7 = $35,000 U

Solution 3:

Materials price variance for March = (SP - AP) * AQ = ($7 - $6.80) * 175000 = $35,000 F

Solution 4:

If Preble had purchased 186,000 pounds of materials at $6.80 per pound and used 175,000 pounds in production, materials quantity variance for March = (SQ - AQ used) * SP = (34000*5 - 175000) * $7 = $35,000 U

Solution 5:

If Preble had purchased 186,000 pounds of materials at $6.80 per pound and used 175,000 pounds in production, materials price variance for March = (SP - AP) * AQ Purchased = ($7 - $6.80) * 186000 = $37,200 F

Note: I have answered more than required parts as HOMEWORKLIB RULES policy, kindly post separate question for answer of remaining questions.